# Machine Learning GPU Graphics Processing Units

### v1

@@ 2.3 Discussion
The approach presented above effectively quantifies uncertainty in trajectory prediction. Yet, it is important to acknowledge some current limitations and potential solutions. One notable concern is the increased memory and computational burden, which may be prohibitive for a real-time application such as trajectory prediction. A possible avenue to address this limitation is through the use of ensemble distillation. Studies have shown that it is possible to distill an ensemble into a single model, thereby significantly reducing computational overhead while maintaining comparable accuracy Malinin et al. (2019). These technique offers a promising direction for future work, ensuring that our approach remains both efficient and performant.
3 Experiments
@@
In this paper, we introduce a novel information-theoretic approach to measure and decompose the uncertainty of the predictive distribution of trajectory prediction models in the AD domain. We model the approximate posterior $q({\mathcal{W}})$ over neural network weights via sampling-based methods, such as dropout (Gal & Ghahramani, 2016) and deep ensembles Lakshminarayanan et al. (2017). For simplicity, we refer to any collection of neural networks as an ensemble. Our experimental analysis is divided into four parts, where we explore both the uncertainty quantification capabilities of our method and the impact of different ensemble compositions. First, in Sec. 3.1, we benchmark our method against an alternative approach to quantify the uncertainty on the original nuScenes dataset (Caesar et al., 2020), which is a commonly used real-world trajectory prediction dataset for AD. We measure the correlation between the uncertainty and the prediction error and explore how epistemic and aleatoric uncertainties complement each other. In the subsequent parts, we create artificial OOD scenarios by manipulating the nuScenes dataset in various ways. Specifically, we propose four different methods for manipulating the original nuScenes dataset, such as removing lane information or omitting parts of the past trajectories of various agents. A detailed explanation of our nuScenes manipulations is provided in App. D. In the second experimental part in Sec. 3.2, we examine the robustness of various models and ensembles across different OOD scenarios. We observe an overall increase in prediction error, indicating that our artificial OOD scenarios are more challenging than the original dataset. In the third part in Sec. 3.3, we investigate how the correlation between uncertainty and prediction error is affected in these OOD scenarios. Lastly, in Sec. 3.4, we study whether we can detect OOD scenarios by analyzing the different types of uncertainty.
@@
Throughout our experiments, we use our novel method to measure the total uncertainty and decompose it into aleatoric and epistemic components to understand their relative importance. We generate trajectory predictions from the ensemble using the approach described in Distelzweig et al. (2024), which involves Model-Based Risk Minimization (MBRM) to draw trajectories from an ensemble of prediction models. For single models, we generate trajectories via Topk sampling, which selects the most likely trajectories (Liu et al., 2024). We rely on LAformer Liu et al. (2024), PGP Deo et al. (2022), and LaPred Kim et al. (2021) to construct different ensembles of trajectory prediction models. These three models are among the best-performing models with available open-source implementations. In our experiments, we evaluate different ensemble configurations, including deep ensembles, dropout ensembles, and single models. We use an ensemble size of three in all experiments; for deep ensembles, we sample three different models, and for dropout ensembles, we sample three different masks. Prediction performance is assessed in terms of minADE and Minimum Final Displacement Error (minFDE) over $k$ proposals. The $\text{minADE}_{k}$ measures the average point-wise L2 distances between the predicted trajectories and the ground truth, returning the minimum over the $k$ proposals (Caesar et al., 2020). In contrast, $\text{minFDE}_{k}$ considers only the final predicted point. Detailed prediction results for different models and ensemble configurations on the original nuScenes dataset are provided in App. C.
3.1 Correlation between Prediction Error and Different Uncertainty Types
@@
Identifying scenarios with high prediction error is critical for safety, as it helps determine when to trust the system or when the driver needs to take control. In this experiment, we study the correlation between different types of uncertainty and prediction error using the original nuScenes dataset. More concretely, we compute the Pearson correlation coefficient $\rho$ between each type of uncertainty and the $\text{minADE}_{k}$. We benchmark our proposed method against Filos et al. (2020), which is an uncertainty quantification approach for planning. Prediction and planning are closely related tasks in AD (Hagedorn et al., 2024), and to the best of our knowledge, Filos et al. (2020) is the only other method with an architecture-agnostic approach that addresses uncertainty quantification in these domains. Filos et al. (2020) estimates the uncertainty by calculating the variance of the log-likelihood of future trajectories with respect to the parameters, i.e., $\text{Var}_{q({\mathcal{W}})}[\log p(y|x,{\mathcal{W}})]$. Unlike our method, this approach requires access to the future trajectory $y$. We report the correlation coefficient between different uncertainty types and the $\text{minADE}_{5}$ in Tab. 1. Additionally, results for $\text{minADE}_{1}$ and $\text{minADE}_{10}$ are provided in App. E.
Table 1: Pearson correlation between $\text{minADE}_{5}$ and different uncertainty types on the original nuScenes dataset. We use sampling via MBRM for ensembles and Topk for single models. LP = LaPred (Kim et al., 2021), LF = LAformer (Liu et al., 2024), PGP (Deo et al., 2022), RIP= Robust Imitative Planning (Filos et al., 2020), Dropout (Gal & Ghahramani, 2016).
@@
We first compare the correlation between the $\text{minADE}_{5}$ and different uncertainty types estimated by our method. We observe that for all ensembles except $3\times\text{LP}$, the total uncertainty has an equal or higher correlation with the prediction error than its individual components, i.e. the aleatoric and epistemic uncertainty. This suggests that both uncertainty sources are complementary. When comparing ensembles against single models, we observe that all ensembles outperform the single models, as these models do not account for epistemic uncertainty. More specifically, when comparing deep ensembles against dropout ensembles, we observe that the former offers a higher correlation with prediction error. This indicates that deep ensembles quantify uncertainty more accurately than dropout, which is in line with the literature on uncertainty quantification with deep ensembles (Lakshminarayanan et al., 2017; Durasov et al., 2021). Lastly, we compare our method against the uncertainty quantification method proposed in Filos et al. (2020), i.e., Robust Imitative Planning (RIP). We observe that our uncertainty quantification method outperforms this approach for all model configurations. This is likely because RIP is based on a heuristic, whereas our method takes a more comprehensive approach. Overall, we observe that the uncertainty estimates obtained by our method provide a useful indication of whether we can trust our model’s predictions or not.
3.2 Robustness of Predictions in OOD Scenarios
@@
In the previous experiment, we analyzed the correlation between uncertainty and prediction error in in-distribution scenarios. We now shift our focus to examining whether prediction performance degrades in OOD scenarios and to what extent. We report the changes in the $\text{minADE}_{5}$ metric with respect to the original dataset in Fig. 3. App. I and App. J provide additional visualizations for $\text{minADE}_{1}$ and $\text{minADE}_{10}$ as well as numerical values.
@@
Overall, we observe that prediction error increases across all datasets in OOD scenarios. However, model ensembles consistently outperform individual models as in all model configurations, more than 50% of the data points fall within the upper green triangle on Fig. 3. This suggests that ensembles offer greater robustness and resilience in OOD scenarios. When comparing deep ensembles composed of the same model to their dropout-based alternatives, the performance is similar when considering the fraction of data points in the green triangle. For instance, the dropout ensemble performs better for PGP, while for LaPred, both the dropout and deep ensemble configurations show equal performance. In contrast, LAformer shows better performance with deep ensembles. However, when examining the mixed deep ensemble that combines different models, we observe a significant increase in performance with all data points lying within the green triangle.
@@
Figure 3: Differences ($\Delta$) in $\text{MinADE}_{5}$ between the original dataset and the corresponding out-of-distribution dataset for baseline models (y-axis) and ensembles (x-axis). Different colors correspond to various baseline models, while different markers denote distinct datasets. Markers positioned in the red area (lower triangle) of each plot indicate that the ensemble exhibits a larger $\Delta\text{MinADE}_{5}$ compared to the baseline. Conversely, markers in the green area signify a smaller $\Delta\text{MinADE}_{5}$ for the ensemble. Percentages indicate how often the ensemble outperforms the baseline.
3.3 Quantifying the Uncertainty in OOD Scenarios
@@
In Sec. 3.1, we investigated whether the uncertainty estimates from our method offer indications of the reliability of our model’s predictions. However, it remains unclear if these findings are also applicable to OOD scenarios. In this experiment, we analyze the correlation between uncertainty and prediction error in OOD scenarios across different ensembles, and we compare these correlation coefficients with those obtained from the original dataset. We report the correlation coefficient between the total uncertainty and the $\text{minADE}_{5}$ in Fig. 4. App. G and App. H provide additional visualizations for $\text{minADE}_{1}$ and $\text{minADE}_{10}$ as well as numerical results.
@@
We first compare the correlation values from the original dataset represented by the circle marker in Fig. 4 with those from the OOD datasets represented by all other markers. The results present a mixed picture – in some OOD scenarios, the correlation coefficient decreases while in others, it increases. Nevertheless, there is a general trend toward a decrease in the correlation coefficient in most OOD cases. Interesting observations can be made when focusing on a specific model ensemble, e.g. the mixed ensemble consisting of LAformer, PGP, and LaPred. Despite being applied to OOD scenarios, this ensemble maintains a higher correlation coefficient than that achieved by the alternative uncertainty quantification method proposed by Filos et al. (2020) on the in-distribution scenarios of the original dataset. This suggests that our proposed method and selected ensemble models offer more robust uncertainty quantification even under challenging conditions.
@@
Next, we investigate whether using an ensemble of models is more advantageous than relying on a single model in OOD scenarios. To assess this, we compare the correlation between uncertainty and prediction error for ensembles versus individual models. Our findings reveal that the ensemble configurations consistently outperform the single model baselines. This conclusion is supported by the fact that in every configuration, more than 50% of the data points lie within the green triangle. This suggests that ensembles provide a more reliable measure of uncertainty in OOD scenarios compared to single models. Lastly, we compare different model ensembles. Specifically, we compare deep ensembles composed of the same model versus their dropout-based alternatives. In two out of three cases, dropout ensembles outperform deep ensembles in terms of the number of data points within the green triangle. However, when we consider a mixed deep ensemble, which combines different models rather than ensembling multiple instances of the same model, we observe a notable improvement in performance. The number of data points within the green triangle increases, signifying that the mixed ensemble achieves a higher correlation between uncertainty and prediction error. This suggests that mixed ensembles, which benefit from a high model diversity, provide better uncertainty quantification than deep ensembles composed of a single model type. This conclusion aligns with our earlier findings, where mixed ensembles consistently performed the best or matched other ensemble configurations in terms of robustness in OOD scenarios. Therefore, we conclude that mixed deep ensembles are the most effective choice for handling OOD scenarios.
@@
Figure 4: Pearson correlation coefficient $\rho$ between total uncertainty and $\text{MinADE}_{5}$ for baseline models ($y$-axis) and ensembles ($x$-axis) over the validation set. Different colors represent various baseline models, while different markers indicate distinct datasets. Markers located in the red area (upper triangle) of each plot signify that the ensemble shows a lower correlation $\rho_{total}$ compared to the baseline. Conversely, markers in the green area (lower triangle) indicate a higher correlation for the ensemble. The numerical value in the bottom right corner of each plot represents the fraction of data points that fall within the green area.
3.4 Detecting OOD Scenarios
@@
In this experiment, our objective is to determine whether we can identify OOD scenarios in the first place. Recognizing such scenarios is critical for improving the performance and robustness of an AD system over time, as it facilitates the collection of challenging cases for re-training and evaluation. We present the uncertainty values for different types of uncertainty in Fig. 5 for both the original nuScenes dataset and various OOD scenarios. For this analysis, we restrict our focus to a mixed deep ensemble consisting of LAformer, PGP, and LaPred, as this ensemble demonstrated the best correlation between uncertainty and prediction error in previous experiments. Additionally, we include OOD detection results for other model ensembles in App. F.
@@
When analyzing epistemic uncertainty, we observe that OOD scenarios exhibit a higher median value than the upper quartile of the original dataset, with the exception of the blackout scenario, where only the median of the original dataset is exceeded. In terms of aleatoric uncertainty, the median for OOD scenarios consistently exceeds the median observed in the original dataset. The total uncertainty follows a similar pattern to aleatoric uncertainty but exhibits a more pronounced difference between OOD and in-distribution cases. These trends indicate that OOD scenarios can be identified with the highest confidence by assessing epistemic uncertainty, a finding that aligns with existing research in uncertainty quantification (Hüllermeier, 2021).
4 Related Work
Anticipating the future motion of traffic participants is a critical component of autonomous driving systems (Hu et al., 2023). Due to the safety-critical nature of these systems, it’s essential to account for and propagate uncertainties across the entire prediction stack (McAllister et al., 2017). For instance, planners need to factor in motion prediction uncertainty to accurately assess the risks associated with various driving maneuvers (Filos et al., 2020). In the following subsections, we review related work on both motion prediction as well as quantification and decomposition of uncertainty.
4.1 Motion Prediction for Autonomous Driving
The future motion of other traffic participants is influenced by a multitude of observable and unobservable factors, rendering it a challenging modeling task. These factors include, among others, the latent goals and preferences of traffic participants, social norms and traffic rules, complex interactions with surrounding traffic, as well as constraints induced by the static environment (Rudenko et al., 2020). The shortcomings of the perception system, which provides noisy and partial observations, pose an additional challenge. These challenges necessitate a probabilistic formulation of the prediction task to adequately model the uncertain and multi-modal nature of future motion. In general, prediction models consist of two components: a behavior backbone, which encodes the traffic scene, and a decoder, which models the predictive distribution. We will highlight various implementations of the two components below.
Early prediction approaches (Casas et al., 2018; Chai et al., 2020; Phan-Minh et al., 2020) propose encoding the past trajectory of observed traffic participants and the elements of the static environment (e.g., lane boundaries, crosswalks, traffic signs) by rendering the scene in a semantic bird’s eye view image and applying well-established convolutional neural networks (He et al., 2016). Such image-based representations of the scene have largely been replaced by vectorized representations due to their inefficiency (Gao et al., 2020; Zhao et al., 2021; Kim et al., 2021; Deo et al., 2022; Nayakanti et al., 2023). In a vectorized representation, all entities of the static and dynamic environment are approximated by a sequence of vectors. Models for sequential data, such as temporal convolutional networks (van den Oord et al., 2016) or recurrent neural networks (Hochreiter & Schmidhuber, 1997; Chung et al., 2014) are used to encode the sequences and interactions between entities are modeled using pooling operations (Alahi et al., 2016), graph neural networks (Hamilton, 2020), or Transformers (Vaswani et al., 2017).
The future motion of traffic participants is typically characterized by a sequence of states over multiple time steps, known as trajectories (Ngiam et al., 2022; Varadarajan et al., 2022). Several strategies are employed to capture the highly multi-modal distribution over trajectories conditioned on the encoded scene. Many approaches represent the distribution by a set of trajectories with associated mode probabilities. The trajectories are either regressed by the model (Cui et al., 2019; Liang et al., 2020; Kim et al., 2021; Deo et al., 2022) or fixed a priori (Phan-Minh et al., 2020). Other approaches use parametric mixture distributions, such as GMMs (Khandelwal et al., 2020; Tolstaya et al., 2021; Varadarajan et al., 2022) or mixtures of Laplacians (Liu et al., 2024). Alternatively, generative models such as conditional variational autoencoders (Lee et al., 2017; Bhattacharyya et al., 2019; Salzmann et al., 2020; Janjoš et al., 2023a), generative adversarial networks (Gupta et al., 2018; Huang et al., 2020; Gómez-Huélamo et al., 2022), normalizing flows (Schöller & Knoll, 2021), or diffusion models (Jiang et al., 2023) model the trajectory distribution via latent variables.
4.2 Uncertainty Modeling, Decomposition and Quantification
The majority of current trajectory prediction models solely account for aleatoric uncertainty by modeling a probability distribution on the output space (Varadarajan et al., 2022). To incorporate epistemic uncertainty in a theoretically sound manner, one can adopt a Bayesian framework (Kendall & Gal, 2017; Depeweg et al., 2018; Wilson & Izmailov, 2020; Wilson, 2020). A Bayesian neural network assumes a distribution over the network weights instead of a point estimate to account for the lack of knowledge about the data-generating process (Hüllermeier, 2021; Jospin et al., 2022). Since analytically evaluating the posterior distribution over the weights is intractable for modern neural networks, approximate inference techniques such as Variational Inference (VI) or forms of MCMC must be considered (Jospin et al., 2022). Due to its simplicity, MC Dropout, which can be interpreted as an approximate VI method (Gal & Ghahramani, 2016), is used by many perception approaches in AD (Kendall & Gal, 2017; Abdar et al., 2021) and is also employed by Janjoš et al. (2023b) for modeling epistemic uncertainty of a trajectory predictor. Another well-established approach to account for epistemic uncertainty is deep ensembles (Lakshminarayanan et al., 2017; Jospin et al., 2022; Wilson & Izmailov, 2020). Prior work (Filos et al., 2020) uses deep ensembles to approximate the posterior distribution in their epistemic uncertainty-aware planning method. We apply MC Dropout as well as deep ensembles to approximate the uncertainty over network weights and systematically assess their performance in the context of trajectory prediction.
A common information-theoretical measure for the uncertainty is the entropy of the predictive distribution as a measure of the total uncertainty, which can be additively decomposed into the conditional entropy and mutual information, representing a measure of aleatoric and epistemic uncertainty (Depeweg et al., 2018; Smith & Gal, 2018; Hüllermeier, 2021; Wimmer et al., 2023). Alternative measures based on the variance were proposed by Depeweg et al. (2018). While variance-based measures are suitable in cases where the predictive distribution is a uni-modal Gaussian, it is less suitable for multi-modal outputs, such as trajectories. Our approach thus relies on entropy-based measures to quantify the uncertainties of trajectory prediction models. However, variance can be useful in other contexts; e.g., Gilles et al. (2022) uses the variance of the predicted heat map over future positions as an uncertainty measure. Another variance-based uncertainty heuristic is proposed by Filos et al. (2020) in the related field of motion planning for AD. This approach, however, only quantifies the epistemic uncertainty and requires access to the future trajectory. Some methods train separate models (Pustynnikov & Eremeev, 2021) or include additional heads with auxiliary tasks (Wiederer et al., 2023; Janjoš et al., 2023b) to learn a proxy measure of the uncertainty of a trajectory prediction model without a proper decomposition.
To the best of our knowledge, we are the first to thoroughly investigate the theoretically sound modeling, decomposition, and quantification of uncertainties for trajectory prediction models. All other published trajectory prediction approaches either do not holistically consider all three aspects or resort to heuristics.
5 Conclusion

### v3

@@ III-C Discussion
The proposed approach effectively quantifies uncertainty in trajectory prediction. However, it is important to acknowledge several current limitations and potential solutions. One notable challenge is the burden of increased memory and computation, which may be prohibitive for real-time applications such as trajectory prediction. A potential solution to this limitation is ensemble distillation, which combines an ensemble of models into a single, more efficient model, significantly reducing computational overhead while maintaining comparable accuracy [54]. A distillation approach for motion prediction models has been proposed in [55]. Alternatively, ensembles can be constructed by modifying only the final layer [56], further mitigating computational costs. These techniques offer a promising direction for future work, ensuring that the approach remains both efficient and performant.
IV Experiments
@@
In this paper, we introduce a novel information-theoretic approach to measure and decompose the uncertainty of the predictive distribution of trajectory prediction models in the AD domain. We model the approximate posterior $q({\mathcal{W}})$ over neural network weights via sampling-based methods, such as dropout [16] and deep ensembles [18]. For simplicity, we refer to any collection of neural networks as an ensemble. Our experimental analysis is divided into four parts, where we explore both the uncertainty quantification capabilities of our method and the impact of different ensemble compositions. First, in Sec. IV-A, we benchmark our method against an alternative approach to quantify the uncertainty on the original nuScenes dataset [31], which is a commonly used real-world trajectory prediction dataset for AD. We measure the correlation between the uncertainty and the prediction error and explore how epistemic and aleatoric uncertainties complement each other. In the subsequent parts, we create artificial OOD scenarios by manipulating the nuScenes dataset in various ways. Specifically, we propose four different methods for manipulating the original nuScenes dataset as described below:
  • RevertEGO: Revert the history of the target vehicle.
  • ScrambleEGO: Randomly shuffle the history of the target vehicle.
  • Blackout: Set 1/2 of the history to zero for the target and all surrounding vehicles.
  • LaneDeletion: Randomly delete 3/4 of all lanes.
Beyond that, we consider combinations of manipulations. In the second experimental part in Sec. IV-B, we examine the robustness of various models and ensembles across different OOD scenarios. We observe an overall increase in prediction error, indicating that our artificial OOD scenarios are more challenging than the original dataset. In the third part in Sec. IV-C, we investigate how the correlation between uncertainty and prediction error is affected in these OOD scenarios. Lastly, in Sec. IV-D, we study whether we can detect OOD scenarios by analyzing the different types of uncertainty.
@@
Throughout our experiments, we use our novel method to measure the total uncertainty and decompose it into aleatoric and epistemic components to understand their relative importance. We generate trajectory predictions from the ensemble using the approach described in [57], which involves Model-Based Risk Minimization (MBRM) to draw trajectories from an ensemble of prediction models. For single models, we generate trajectories via Topk sampling, which selects the most likely trajectories [8]. We rely on LAformer [8], PGP [5], and LaPred [4] to construct different ensembles of trajectory prediction models. These three models are among the best-performing models with available open-source implementations. In our experiments, we evaluate different ensemble configurations, including deep ensembles, dropout ensembles, and single models. We use an ensemble size of three in all experiments; for deep ensembles, we sample three different models, and for dropout ensembles, we sample three different masks. Prediction performance is assessed in terms of Minimum Average Displacement Error (minADE) over $5$ proposals. The $\text{minADE}_{5}$ measures the average point-wise L2 distances between the predicted trajectories and the ground truth, returning the minimum over the proposals [31].
IV-A Correlation between Prediction Error and Different Uncertainty Types
@@
Determining whether predictions can be trusted is crucial for deciding when to rely on the system or when the driver should take control. In this experiment, we analyze the correlation between different types of uncertainty and prediction error using the original nuScenes dataset. More concretely, we compute the Pearson correlation coefficient $\rho$ between each type of uncertainty and the $\text{minADE}_{5}$. We benchmark our proposed method against [32], which is an uncertainty quantification approach for planning. To the best of our knowledge, it is the only other architecture-agnostic method that addresses uncertainty quantification in the domain of autonomous driving. More concretely, [32] estimates uncertainty by computing the variance of the log-likelihood of future trajectories with respect to the parameters, i.e., $\text{Var}_{q({\mathcal{W}})}[\log p(y|x,{\mathcal{W}})]$. We report the $\text{minADE}_{5}$ values along with the correlation coefficient between different uncertainty types and the $\text{minADE}_{5}$ in Tab. I.
Table I: Correlation between Different Uncertainty Types and Prediction Error on the nuScenes Dataset.
@@
$\text{minADE}_{5}$ and Pearson correlation (higher is better) between $\text{minADE}_{5}$ and different uncertainty types on the original nuScenes dataset. We use sampling via MBRM [57] for ensembles and Topk for single models. LP = LaPred [4], LF = LAformer [8], PGP [5], Dropout [16], RIP = Robust Imitative Planning [32].
@@
We first compare the correlation between the $\text{minADE}_{5}$ and different uncertainty types estimated by our method. We observe that for all ensembles except $3\times\text{LP}$, the total uncertainty has an equal or higher correlation with the prediction error than its individual components, i.e. the aleatoric and epistemic uncertainty. This suggests that both uncertainty sources are complementary. When comparing ensembles against single models, we observe that all ensembles outperform the single models, as these models do not account for epistemic uncertainty. Moreover, when comparing deep ensembles against dropout ensembles, we observe that the former offers a higher correlation coefficient. This indicates that deep ensembles quantify uncertainty more accurately than dropout, which is in line with the literature on uncertainty quantification with deep ensembles [18, 58]. Lastly, we compare our method against the uncertainty quantification method proposed in [32], i.e., Robust Imitative Planning (RIP). We observe that our uncertainty quantification method outperforms this approach for all model configurations. This is likely because RIP is based on a heuristic, whereas our method takes a more comprehensive approach. Overall, we observe that the uncertainty estimates obtained by our method provide a useful indication of whether we can trust our model’s predictions or not.
IV-B Robustness of Predictions in OOD Scenarios
@@
In the previous experiment, we analyzed the correlation between uncertainty and prediction error in In-Distribution (ID) scenarios. We now shift our focus to examining whether prediction performance degrades in OOD scenarios and to what extent. We report the changes in the $\text{minADE}_{5}$ metric with respect to the original dataset in Fig. 3.
@@
Overall, we observe that prediction error increases across all datasets in OOD scenarios, indicating that our dataset augmentations create a more challenging evaluation setting. However, model ensembles consistently outperform individual models, as more than 50% of the data points fall within the upper green triangle in Fig. 3 across all model configurations. This suggests that ensembles provide greater robustness and resilience in OOD scenarios. When comparing deep ensembles composed of the same model to their dropout-based counterparts, performance remains similar in terms of the fraction of data points in the green triangle. For example, the dropout ensemble outperforms deep ensembles for PGP, while LaPred exhibits equal performance across both configurations. In contrast, LAformer benefits more from deep ensembles. Notably, when evaluating the mixed deep ensemble, which combines different models, we observe a substantial performance improvement, with all data points falling within the green triangle.
@@
Figure 3: Differences ($\Delta$) in $\text{MinADE}_{5}$ between the original dataset and the corresponding out-of-distribution dataset for baseline models (y-axis) and ensembles (x-axis). Different colors correspond to various baseline models, while different markers denote distinct dataset augmentations. Markers positioned in the red area (lower triangle) of each plot indicate that the ensemble exhibits a larger $\Delta\text{MinADE}_{5}$ compared to the baseline. Conversely, markers in the green area signify a smaller $\Delta\text{MinADE}_{5}$ for the ensemble. Percentages indicate how often the ensemble outperforms the baseline. Upper row represents deep ensembles and lower row Dropout ensembles.
IV-C Quantifying the Uncertainty in OOD Scenarios
@@
In Sec. IV-A, we investigated whether the uncertainty estimates from our method offer indications of the reliability of our model’s predictions. However, it remains unclear if these findings are also applicable to OOD scenarios. Specifically, can we trust our uncertainty estimates when encountering out-of-distribution inputs? In this experiment, we analyze the correlation between uncertainty and prediction error in OOD scenarios across different ensembles, and we compare these correlation coefficients with those obtained from the original dataset. We report the correlation coefficient between the total uncertainty and the $\text{minADE}_{5}$ in Fig. 4.
@@
We first compare the correlation values from the original dataset represented by the circle marker in Fig. 4 with those from the OOD datasets represented by all other markers. The results present a mixed picture – in some OOD scenarios, the correlation coefficient decreases while in others, it increases. Nevertheless, there is a general trend toward a decrease in the correlation coefficient in most OOD cases.
Interesting insights emerge when comparing the results of our approach with the results of the RIP uncertainty quantification approach [32]. Since RIP estimates only epistemic uncertainty, we evaluate the Pearson correlation coefficient between epistemic uncertainty and $\text{minADE}_{5}$ across all OOD scenarios. Due to space constraints, we provide only a summary of the results. In 41 out of 42 examined ensemble configuration and OOD scenario combinations, our approach yields a higher correlation coefficient than RIP. Notably, the average correlation increase is most pronounced in the mixed ensemble and LaPred ensemble, with improvements of 406% and 434% over RIP, respectively. The smallest average increase is observed in the dropout PGP ensemble configuration, at 61%. These results suggest that our method provides more robust uncertainty quantification, even in challenging OOD conditions.
Next, we investigate whether using an ensemble of models is more beneficial than relying on a single model in OOD scenarios. To evaluate this, we compare the correlation between uncertainty and prediction error for ensembles versus individual models in Fig. 4. Our results consistently show that ensemble configurations outperform single-model baselines. This conclusion is reinforced by the fact that in every configuration, more than 50% of the data points lie within the green triangle, indicating that ensembles provide a more reliable measure of uncertainty in OOD scenarios compared to individual models.
@@
We then compare three different ensemble configurations in Fig. 4, specifically (i) dropout-based ensembles, (ii) deep ensembles composed of the same model, and (iii) mixed deep ensembles composed of Laformer, PGP, and LaPred. In two out of three cases, (i) outperforms (ii) in terms of the number of markers within the green triangle. However, when considering (iii), we observe that its markers are fully in the green triangle. This is a notable performance improvement compared to (ii) as well as (i), which manages to match the mixed ensembles only in a single configuration. These findings suggest that mixed ensembles, which benefit from increased model diversity, provide superior uncertainty quantification compared to other methods. This conclusion aligns with our previous results in Fig. 3, where mixed ensembles consistently performed the best or matched other ensemble configurations in terms of robustness in OOD scenarios. Therefore, we conclude that mixed deep ensembles are the most effective choice for handling OOD scenarios.
@@
Figure 4: Pearson correlation coefficient $\rho$ between total uncertainty and $\text{MinADE}_{5}$ for baseline models ($y$-axis) and ensembles ($x$-axis) over the validation set. Different colors represent various baseline models, while different markers indicate distinct dataset augmentations. Markers located in the red area (upper triangle) of each plot signify that the ensemble shows a lower correlation $\rho_{total}$ compared to the baseline. Conversely, markers in the green area (lower triangle) indicate a higher correlation for the ensemble. The numerical value in the bottom right corner of each plot represents the fraction of data points that fall within the green area. Upper row represents deep ensembles and lower row Dropout ensembles.
IV-D Detecting OOD Scenarios
@@
In this experiment, our objective is to determine whether OOD scenarios can be reliably identified. Detecting such scenarios is crucial for safety, as it enables the autonomous system to alert the driver when intervention is necessary. Additionally, recognizing OOD cases enhances the performance and robustness of an AD system over time by facilitating the collection of challenging instances for re-training and evaluation. We present the uncertainty values for different types of uncertainty in Fig. 5 for both the original nuScenes dataset and various OOD scenarios. For this analysis, we restrict our focus to a mixed deep ensemble consisting of LAformer, PGP, and LaPred, as this ensemble was favorable in terms of calibrations and robustness in previous experiments.
@@
When analyzing epistemic uncertainty, we observe that OOD scenarios exhibit a higher median value than the upper quartile of the original dataset, with the exception of the blackout scenario, where only the median of the original dataset is exceeded. In terms of aleatoric uncertainty, the median for OOD scenarios consistently exceeds the median observed in the original dataset. The total uncertainty follows a similar pattern to aleatoric uncertainty but exhibits a more pronounced difference between OOD and ID cases. These trends indicate that OOD scenarios can be identified with highest confidence by assessing epistemic uncertainty, a finding that aligns with existing research in uncertainty quantification [27].
V Conclusion